OpenStudy (anonymous):

can someone put -5x+10y=5 into y = for me me ? will give medal.

OpenStudy (anonymous):

y=1/2x+1/2

OpenStudy (aggie12341):

I mean y=1/2x+1/2

OpenStudy (anonymous):

Heyyy, that's bribery! Confusticate! \[y = \frac{ 5x +5 }{ 10 } = \frac{ x+1 }{ 2 }\]

OpenStudy (anonymous):

y=1/2x+1/2

OpenStudy (aggie12341):

medal!

OpenStudy (aggie12341):

Your wrong anyways!

OpenStudy (anonymous):

I'm perfectly right

OpenStudy (anonymous):

I was first but whatever

OpenStudy (aggie12341):

No honey I was first go look. Are you blind?

OpenStudy (anonymous):

THIS ANSWER IS WRONG: \[y = \frac{ 1 }{ 2x }+\frac{ 1 }{ 2 }\]

OpenStudy (aggie12341):

I got a medal so.....:P

OpenStudy (anonymous):

Doesn't alter the fact that you're wrong